yes MATH 270 Probability and Statistics Spring 2018 Homework: Assignment 5 Score: 0 of 1 pt 5.3.18 HW Score: 47.5%, 19 of 40 pts E Question Help A book can be classified as either non-fiction or fiction Suppose that 99% of books are classified as (a) Two books are chosen at rando (b) Six books are chosen at random. What is the probability that all six books are fiction? ction. m What is the probability that both books are fiction? hat at least one of six randomly selected books is non-fiction? Would it be unusual that at least one of sik randomly selected books is non-fiction? (a) Two books are chosen at random. What is the probability that both books are fiction? Round to four decimal places as needed) rch
Answer #1

99% of books are classified as fiction. Probability of a chosen book classified as fiction : P(F) = 99/100 = 0.99 Probability of a chosen book classified as non-fiction P(NF) = 1-0.99 =0.01 (a) the probability is (0.99)2 = 0.9801

(b) the probability is (0.99)6 = 0.9415

(c) 1 - prob that none of the 6 are fraction = 1 - (0.99)6

= 1 - 0.9415 = 0.0585

When probability is less than 5% = 0.05 then these are unusual.

Here probablity is greater than 0.05 then it is usual.